OpenStudy (lena772):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (lena772):

@hartnn

OpenStudy (lena772):

@ranga

OpenStudy (lena772):

@shamil98

OpenStudy (shamil98):

2x + 3y= 1200 change this to slope-intercept form y = mx + b

OpenStudy (anonymous):

The slope-intercept form is y=mx+b where m is the gradient/slope and b is the y-intercept. So if you rearrange the equation... 2x + 3y = 1200 You want only y on the left side. So you will have to get rid of the 2x and scalar 3 on the left side. To do that, you will minus both sides with 2x to get 2x + 3y -2x = 1200 -2x 2y = 1200 - 2x Now you just divide it by 2

OpenStudy (shamil98):

where does the 3y go ? @amonaroll32 2y?...

OpenStudy (shamil98):

it's not 2y = 1200 -2x it's 3y = -2x + 1200 you divide everything by 3 :)

OpenStudy (anonymous):

Shoot, sorry about that. Yeah, @shamil98 is correct :)